OpenStudy (anonymous):

A boy is flying a kite from a string of length 150m. If the string is taut and makes an angle of 67 degrees with the horizontal, what is the height of the kite?

OpenStudy (tkhunny):

\(\sin(67º) = \dfrac{Height}{150\;m}\) This should look familiar. There is a Right Triangle in there, somewhere.

OpenStudy (anonymous):

answer is 67.1 m

OpenStudy (tkhunny):

Try again. \(\sin(67º) = 0.83867\)